OpenStudy (anonymous):

simplify to a single logarithm, using logarithm properties: log(12x^4)-log(4x)

OpenStudy (anonymous):

Log rules state that log(a) - log(b) = log(a/b)

OpenStudy (anonymous):

what do u mean? so 12x^4/4x

OpenStudy (anonymous):

Yep, just simplify\[\log_{?} \frac{ 12x^4 }{ 4x}\] like any fraction

OpenStudy (anonymous):

so: log 3x^4/x

OpenStudy (anonymous):

close, simplify your x's. x^4/ x = x^3

OpenStudy (anonymous):

so the answer would be log3x^3

OpenStudy (anonymous):

That's it!